OpenStudy (anonymous):

lets see who can get this answer -5(n-3)^2=10

OpenStudy (anonymous):

show ur steps

jhonyy9 (jhonyy9):

divide both sides by -5 and will get (n-3)^2 = -2 but you know that (n-3)^2 = n^2 -6n +9 so n^2 -6n +9= -2 n^2 -6n +11 =0 so now just you need to solve this quadratic for n